OpenStudy (anonymous):

please help? 'find the next three terms of the arithmetic sequence: 39, 50, 61, 72 ...'

OpenStudy (anonymous):

Add 11

OpenStudy (anonymous):

72 + 11= 83 83 + 11= 94 94+ 11= 105

OpenStudy (anonymous):

haha. I figured it out. srry.. I wasn't thinking stright..

OpenStudy (the_fizicx99):

Find the difference, 61-50 = 11 72-61 = 11 And so on, the difference is 11, formula is a_n = a_1 + d(n - 1) a_5 = 39 + 11(5 - 1) a_5 = 39 + 11(4) a_5 = 39 + 44 a_5 = 83 So the fifth term is 83, you can continue doing so for the remaining 2 digits if you like.

OpenStudy (anonymous):

Using the Table function from Mathematica: Table[ 39+11(n-1), {n, 7} ] -> {39, 50, 61, 72, 83, 94, 105}